# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmarking of Large Language Models in Mental Health Question Answering

### v2

@@ 4.6 Can LLMs Reliably Judge the Quality of Responses?
Human expert evaluation provides gold-standard insight but is costly and difficult to scale (Gehrmann et al., 2021). As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality (Madaan et al., 2023; Manakul et al., 2023), but its reliability in high-stakes, subjective domains like mental health remains unclear. To investigate this, we tested eight advanced LLMs as automated judges, as listed in Figure 3. Each model was prompted with the same QA pairs used in Section 4.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model (see Appendix LABEL:sub_appendeix_llm_judge_eval), and additional results for LLM judges are in Appendix LABEL:llm_judge_additional.
Figure 3: Average evaluation scores across six dimensions (subplots) for counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including nine LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
@@
LLM Rankings Diverge from Human Preferences. Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 3 and Figure 3. The preferences expressed by LLM judges diverged sharply from those of human annotators, except for GPT-5 (dark blue line in Table 3), which calls into question the reliability of LLMs in identifying relative quality. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge for Overall Quality.
@@ 5 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures
Identifying Fine-grained Failure Modes To construct a targeted adversarial benchmark, we first need to isolate that LLMs exhibit. Broad categories (e.g., “lacking personalization”) from Section 4.5 are too general. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract concrete failure modes:
@@
Adversarial Dataset Construction We rehired 10 mental health professionals who previously participated in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic mental health questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error. Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated responses.
@@
Evaluation Setup. We used the same prompting setup as in CounselBench-EVAL for these adversarial questions with nine LLMs (results are in Table 3). For each model, we issued the same prompt three times to sample variations in generation. In total, this yielded 3,240 QA pairs (120 questions × 9 models × 3 generations). To determine whether a response triggered its targeted failure mode, we used GPT-4.1 as a scalable labeling tool. The task was deliberately framed as binary classification with clear clinician-authored definitions and one in-context example per failure type (§LABEL:sec:aversarial_by_gpt_41), making it simpler and more reproducible than multi-dimensional quality scoring.
Human Validation. A mental health professional spot-checked 96 responses, yielding 72.9% agreement with GPT-4.1 (Stemler, 2004). This level of concordance indicates that GPT-4.1 can serve as a practical proxy for detecting whether a predefined failure was triggered, though it is not a substitute for full expert evaluation. We therefore use GPT-based labels only to quantify trigger rates in CounselBench-Adv, while all primary analyses of quality and safety remain grounded in expert annotations from CounselBench-Eval. Expanding human annotation to cover the full set of 3,240 adversarial responses represents an important direction for future work.
@@
Table 3: Fraction of responses identified to contain each targeted failure mode. Higher values reflect greater model vulnerability to the targeted issue. Red: zero-shot; Orange: few-shot.
@@
Specific Issue GPT-3.5-Turbo GPT-4 GPT-5 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.25   0.32   0.97   0.40   0.55   0.47   0.53   0.42   0.47 2. Therapy   0.28   0.25   0.93   0.60   0.53   0.50   0.53   0.30   0.35 3. Symptoms   0.20   0.82   0.85   0.83   0.83   0.87   0.83   0.67   0.68 4. Judgmental   0.12   0.02   0.02   0.22   0.22   0.15   0.08   0.15   0.08 5. Apathetic   0.85   0.22 0   0.18   0.13   0.10   0.12   0.47   0.42 6. Assumptions   0.53   0.35   0.17   0.73   0.80   0.57   0.32   0.42   0.48 1. Medication   0.25   0.47   0.92   0.45   0.48   0.62   0.62   0.50   0.57 2. Therapy   0.43   0.40   0.95   0.78   0.75   0.65   0.58   0.40   0.50 3. Symptoms   0.42   0.83   0.85   0.82   0.82   0.85   0.82   0.72   0.65 4. Judgmental   0.08   0.08   0.07   0.22   0.20   0.15   0.17   0.05   0.15 5. Apathetic   0.42   0.07 0   0.17   0.07   0.08   0.03   0.12   0.50 6. Assumptions   0.45   0.67   0.15   0.80   0.82   0.67   0.55   0.42   0.65
Adversarial Questions Effectively Trigger Targeted Failures Table 3 reports the frequency of each failure mode across model outputs based on GPT-4.1 evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Notably, speculation about medical symptoms emerged as the most prevalent failure: it was triggered in 82% of GPT-4 responses, 83% of LLaMA series responses, over 83% of Claude models, and more than 67% of Gemini outputs. This pattern raises serious concerns, as such speculation can lead to inappropriate clinical guidance without proper diagnosis. In contrast, judgmental or apathetic tone was the less common issues across models.
To further explore potential causes and mitigation strategies, we conducted a follow-up experiment on CounselBench-ADV using few-shot prompting (details described in Appendix LABEL:sec:few-shot-adv). The resulting failure rates are also shown in Table 3. Across most categories, few-shot prompting yielded limited improvements. This suggests that many failure modes are likely rooted in pretraining and not easily resolved through in-context examples.
@@ 6 Discussion
This work presents CounselBench, a benchmark that evaluates LLMs on open-ended mental health QA. Unlike prior resources that rely on multiple-choice tasks or small-scale expert review, it brings together two complementary components: a large-scale corpus of 2,000 expert evaluations with detailed written rationales (CounselBench-Eval), and a set of 120 adversarial questions written by clinicians to expose failure modes (CounselBench-Adv). Together, they provide both breadth and depth: broad coverage through systematic expert scoring, and targeted probing through adversarial design. Our analysis shows that LLMs can achieve high ratings on several dimensions of quality, yet continue to exhibit recurring weaknesses, including unconstructive feedback, overgeneralization, limited personalization, and unauthorized medical advice. These findings highlight both the progress and the open challenges in deploying LLMs responsibly in mental health QA.
@@
Both components of CounselBench are released as public resources to support future research. For example, CounselBench-Eval can be used to (1) train and validate alignment methods such as reinforcement learning from human feedback, (2) develop automated quality or safety detectors grounded in span-level annotations, and (3) study expert rationales to build critique models or interpretability tools. CounselBench-Adv enables complementary research, such as (1) stress-testing new models to compare robustness under clinically relevant failure modes, (2) incorporating adversarial cases into red-teaming and auditing pipelines, and (3) using these targeted examples as hard negatives for fine-tuning or as feedback signals in multi-agent critique systems. By releasing these resources, we aim to provide a foundation for both rigorous benchmarking and practical model improvement in mental health QA and related high-stakes applications.
@@ Acknowledgments
We release all data and source code. Full hyperparameters, implementation details, random seeds, and prompt templates are in Appendix D, and computational resources are detailed in Appendix LABEL:compute_resouces, to ensure replicability.
@@ Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 4 and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers — one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an human therapist. For each post-response pair, annotators evaluated across our metrics, providing rationals and span-level annotations. At the end of the survey, we collected demographic and professional background data, including gender, ethnicity, freelance capacity, counseling experience, credentials, and areas of specialization.
Figure 4: Survey interface: annotators read a user post and one response (left) and rate the response on criteria with Likert scales and text‐evidence boxes (right).

### v3

@@ 4.6 Can LLMs Reliably Judge the Quality of Responses?
Human expert evaluation provides gold-standard insight but is costly and difficult to scale (Gehrmann et al., 2021). As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality (Madaan et al., 2023; Manakul et al., 2023), but its reliability in high-stakes, subjective domains like mental health remains unclear. To investigate this, we tested nine advanced LLMs as automated judges, as listed in Figure 3. Each model was prompted with the same QA pairs used in Section 4.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model (see Appendix LABEL:sub_appendeix_llm_judge_eval), and additional results for LLM judges are in Appendix LABEL:llm_judge_additional.
Figure 3: Average evaluation scores across six dimensions (subplots) for responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including nine LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
@@
LLM Rankings Diverge from Human Preferences. Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 3. The preferences expressed by LLM judges diverged sharply from those of human annotators, except for GPT-5 (dark blue line in Table 3), which calls into question the reliability of LLMs in identifying relative quality. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge for Overall Quality.
@@ 5 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures
Identifying Fine-grained Failure Modes. To construct a targeted adversarial benchmark, we first need to isolate the failure modes that LLMs exhibit. Broad categories (e.g., “lacking personalization”) from Section 4.5 are too general. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract concrete failure modes:
@@
Adversarial Dataset Construction. We rehired 10 mental health professionals who previously participated in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic mental health questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error (see the questionnaire screenshot in Appendix 4). Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated responses.
@@
Evaluation Setup. We used the same prompting setup as in CounselBench-EVAL for these adversarial questions with nine LLMs. For each model, we generated one response for each question, this leads to 1,080 QA pairs (120 questions × 9 LLMs). The annotation task was deliberately framed as classification (“yes”/ “no”/“not sure”) with clear clinician-authored definitions and one in-context example per failure type (Appendix LABEL:appendix_failure_modes), rather than the multi-dimensional metrics used in CounselBench-Eval. Five mental-health practitioners, different from those who authored the adversarial questions, then annotated whether each response has the targeted issues. The full annotation setup is described in Appendix LABEL:sec:adv_llm_judge_results.
@@
Table 3: Fraction of responses identified to contain each targeted failure mode by five mental health professionals. Higher values reflect greater model vulnerability to the targeted issue.
@@
Specific Issue GPT-3.5-Turbo GPT-4 GPT-5 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.05 0   0.47   0.05   0.1 0 0 0 0 2. Therapy   0.2   0.2   0.85   0.55   0.65   0.45   0.5   0.2   0.26 3. Symptoms   0.15   0.45   0.6   0.45   0.45   0.5   0.37   0.26   0.25 4. Judgmental   0.25   0.25   0.05   0.11   0.1   0.05   0.1   0.2   0.1 5. Apathetic   0.7   0.2   0.15   0.15   0.15   0.05   0.2   0.4   0.3 6. Assumptions   0.4   0.35   0.15   0.25   0.25   0.35   0.25   0.4   0.35
Adversarial Questions Effectively Trigger Targeted Failures. Table 3 reports the frequency of each failure mode across model outputs based on expert evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Several patterns stand out: therapy suggestions are highest for GPT‑5 (0.85), substantial for Llama models (0.55–0.65), and present for Claude models (0.45–0.50). Symptoms speculation is also frequent in the three above model families, and lower for Gemini (0.25-0.26) and GPT‑3.5‑Turbo (0.15). Unsupported assumptions appear in most models (0.25–0.40), with GPT‑5 lower at 0.15. Medication advice is rare for nearly all models (0–0.10), but GPT‑5 is an outlier at 0.47. Apathetic tone peaks for GPT‑3.5‑Turbo (0.70), while judgmental tone remains comparatively low across models (0.05–0.25).
@@
Table 4: LLM-as-Judge performance on CounselBench-Adv, relative to human expert ratings.
Metrics LM-as-Judges GPT-3.5-turbo GPT-4 GPT-4.1 GPT-5 Llama-3.1-70B-Instruct Llama-3.3-70B-Instruct Claude-3.7-sonnet Gemini-2.0-flash Acc. 0.74 0.70 0.64 0.67 0.63 0.64 0.70 0.63 F1 0.41 0.35 0.49 0.49 0.48 0.48 0.50 0.46
LLM Judges Remain Misaligned with Human Expert Judgments.
We also evaluated an LLM-as-Judge setup for failure-mode detection across eight models  [^1] Gemini-1.5-Pro and Claude-3.5-Sonnet, which are included in the LLM-as-judge analysis in Section §4.6, were not longer available when we ran the LLM-as-judge experiments on CounselBench-ADV. . The experiment follows the same setup in Section §LABEL:sub_appendeix_llm_judge_eval with LABEL:prompt:gpt4.1-failure-mode, and additional details are provided in Appendix LABEL:sec:adv_llm_judge_results. Despite providing explicit definitions and in-context examples, we observe a substantial gap between LLMs and expert judgements (Table 4). Even the best-performing model (Claude-3.7-Sonnet) attains an F1 score of only 0.5. The performance breakdown of each LLM judge for each failure mode is present in Figure LABEL:fig:tpr_fpr_llm_judge_adv. Additional analysis for the LLM judges is provided in Appendix LABEL:sec:adv_llm_judge_results.
@@ 6 Discussion
This work presents CounselBench, a benchmark that evaluates LLMs on open-ended mental health QA. Unlike prior resources that rely on multiple-choice tasks or small-scale expert review, it brings together two complementary components: a large-scale corpus of 2,000 expert evaluations with detailed written rationales (CounselBench-Eval), and a set of 120 adversarial questions written by clinicians to expose failure modes paired with 1,080 expert annotations for failure mode detection (CounselBench-Adv). Together, they provide both breadth and depth: broad coverage through systematic expert scoring, and targeted probing through adversarial design. Our analysis shows that LLMs can achieve high ratings on several dimensions of quality, yet continue to exhibit recurring weaknesses, including unconstructive feedback, overgeneralization, limited personalization, and unauthorized medical advice. These findings highlight both the progress and the open challenges in deploying LLMs responsibly in mental health QA.
As noted in prior surveys of mental‑health datasets (Harrigian et al., 2021; Mandal et al., 2025), publicly available data with clinician responses are extremely scarce. On one hand, most clinical conversations and patient-provider messages cannot be released due to privacy protections (e.g., de-identification risks highlighted by Kovačević et al. (2024)); on the other hand, peer‑support forums (Marshall et al., 2024) and social‑media platforms such as Reddit (Boettcher, 2021) typically contain non‑clinical, user‑generated content with noisy labels and without professional verification (Garg, 2023). Within this landscape, CounselChat remains one of the few openly accessible resources that pair patient questions with verified therapist responses, and prior work has examined its relatively high quality and utility for research (Ahmed et al., 2024). Although current public datasets are limited, the evaluation framework in CounselBench, including the six-dimension rubric and annotation protocol, can be applied directly to future datasets. This structure can support consistent model comparison across broader settings in subsequent work.
While CounselBench is designed for evaluating open-ended mental health QA, our framework provides a foundation for future expansion into dialogue settings. For CounselBench-Eval, the rubric can be applied turn-by-turn or at the session level (Chen et al., 2025) and aggregated across the dialogue, for example, by averaging scores or weighting them based on conversational role (e.g., first vs. follow-up turns). Moving to multi-turn exchanges, however, introduces additional complexity, including the need for models to appropriately track and utilize prior conversational context. This raises challenges in both prompting design (e.g., how much history to include) and evaluation (e.g., how to measure coherence or consistency over time). For CounselBench-ADV, extending to multi-turn adversarial prompts requires preserving interactional dynamics while targeting specific failure modes. Promising directions include using simulated patient agents (Wang et al., 2024a; b) to scaffold coherent multi-turn interactions that elicit subtle failures, and developing multi-turn red-teaming techniques seeded from our identified adversarial examples (Guo et al., 2025).
@@
Both components of CounselBench are released as public resources to support future research. CounselBench-Eval can be used to (1) train and validate alignment methods such as reinforcement learning from human feedback (Ouyang et al., 2022; Zhang et al., 2024; Li et al., 2025a), (2) develop automated quality or safety detectors grounded in span-level annotations (Zhang et al., 2023), and (3) study expert rationales to build critique models or interpretability tools. CounselBench-Adv enables complementary research, such as (1) stress-testing new models to compare robustness under clinically relevant failure modes, (2) incorporating adversarial cases into red-teaming and auditing pipelines (Mazeika et al., 2024), (3) using these targeted examples as hard negatives for fine-tuning or as feedback signals in multi-agent critique systems (Paulus et al., 2025; Gomaa et al., 2025). By releasing these resources, we aim to support both rigorous benchmarking and practical model improvement in mental health QA and related high-stakes applications.
Ethics Statement
@@ Acknowledgments
We release the data and source code. Full hyperparameters, implementation details, random seeds, and prompt templates are in Appendix D, and computational resources are detailed in Appendix LABEL:compute_resouces, to ensure replicability.
@@ Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 4 for CounelBench-Eval and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers - one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an online human therapist. For each post-response pair, annotators evaluated across our metrics, providing rationals and span-level annotations. At the end of the survey, we collected demographic and professional background data, including gender, ethnicity, freelance capacity, counseling experience, credentials, and areas of specialization.
Figure 5 shows a screenshot for collecting CounselBench-ADV questions. Each annotator was asked to write two new questions per failure mode, which could trigger the targeted failure mode. The failure mode descriptions/examples/details for the questionnaires are listed in TableLABEL:tab:six_issues.
Figure 4: Survey interface CounselBench-Eval: annotators read a user post and one response (left) and rate the response on criteria with Likert scales and text‐evidence boxes (right).
Figure 5: Survey Interface for CounselBench-ADV: annotators are given the failure mode, examples, and explanations for the targeted failure mode and are asked to draft new questions that could trigger the same failure mode.